# The Central Role of the Loss Function in Reinforcement Learning

### v2

@@ 4 Reinforcement Learning
In the following sections, we will see that these observations and insights generally transfer to more complex decision making setups, in particular reinforcement learning (RL). Compared to the CSC setting, two new challenges of RL are that (1) the learner receives feedback only for the chosen action (a.k.a., partial or bandit feedback) and (2) the learner sequentially interacts with the environment over multiple time steps. As before, we focus on value-based algorithms with function approximation and prove bounds for problems with high-dimensional observations, i.e., beyond the finite tabular setting.
@@ 4.1 Problem Setup
4.2 Solving Online RL with Optimistic Squared-Loss Regression
Algorithm 1 Policy Roll-In
@@ 4.2 Solving Online RL with Optimistic Squared-Loss Regression
BC ensures that the TD-style regression which bootstraps on the next prediction is realizable, playing the same role as realizability (Assump. 1) in the CSC setting. In fact, BC implies realizability in $Q^{\star}$: $Q_{h}^{\star}\in\mathcal{F}_{h}$ for all $h$, which can be verified by using the Bellman optimality equations and induction from $h=H\to 1$. While appealing, $Q^{\star}$-realizability is not sufficient for sample efficient RL [59, 26] and TD learning can diverge or converge to bad points with realizability alone [56, 50, 37]. We note that $Q^{\star}$-realizability becomes sufficient when combined with other types of assumptions such as generative access to the MDP [46], where the learner can reset to any previously observed states. We believe that the techniques in this paper can lead to first- and second-order bounds with realizability plus generative access for example. However, we do not pursue this direction here since exchanging BC for other conditions is orthogonal to our study of loss functions.
@@
Fix any set $\mathcal{S}$, function class $\Psi=\{\psi:\mathcal{S}\to\Delta(\mathbb{R})\}$, distribution class =$\mathcal{M}=\{\nu:\Delta(\mathcal{S})\}$, threshold $\varepsilon_{0}$, and number $q\in\mathbb{N}$. The $\ell_{q}$-eluder dimension $\operatorname{EluDim}_{q}(\Psi,\mathcal{P},\varepsilon_{0})$ is the length of the longest sequence $p^{(1)},\dots,p^{(L)}\subset\mathcal{P}$ s.t. $\exists\varepsilon\geq\varepsilon_{0}$, $\forall t\in[L]$, $\exists\psi\in\Psi$ s.t. $\left\lvert\mathbb{E}_{p^{(t)}}\psi\right\rvert>\varepsilon$ but $\sum_{i<t}\left\lvert\mathbb{E}_{p^{(i)}}\psi\right\rvert^{q}\leq\varepsilon^{q}$.
@@
The distribution class $\mathcal{M}$ will be the set of all visitation distributions by any policy, i.e., $\mathcal{M}_{h}=\{d^{\pi}_{h}(\cdot):\pi\in\Pi\}$ where $d^{\pi}_{h}(x,a)$ is the state-action visitation distribution of $\pi$ at time step $h$. If UA is enabled, then we will have $\mathcal{S}=\mathcal{X}$ and $\Psi_{h}^{\operatorname{sq},\operatorname{V}}=\{\mathbb{E}_{a\sim\operatorname{unif}(\mathcal{A})}[\psi(x,a)]:\psi\in\Psi^{\operatorname{sq}}_{h}\}$ simply takes uniform distribution for the action argument, where the $V$ superscript denotes that this is ‘V-type’. The V-type distribution class is the set of state visitation distributions $d^{\pi}_{h}(x)$ at time step $h$. Thus, define the eluder dimension for squared loss:
@@
We now state the guarantees for Alg. 2 with squared loss $\ell_{\operatorname{sq}}$, which recovers the results of [33].
@@
Under Assump. 3, for any $\delta\in(0,1)$, w.p.a.l. $1-\delta$, Alg. 2 with the squared loss $\ell_{\operatorname{sq}}$ and $\beta=2\ln(H|\mathcal{F}|/\delta)$ enjoys the following:
@@
Computational complexity of version space algorithms. While we presented version space algorithms for their simplicity and statistical efficiency, we note here that they are computationally hard to run in general. Specifically, the computational cost of optimizing over the version space is NP-hard even in tabular MDPs [17]. However, in the one-step $H=1$ setting (a.k.a. contextual bandits), the version space optimization is oracle-efficient [23, 20, 62]. In the RL setting, there are also approaches to mitigate the computational hardness of optimism. One approach is to use $\varepsilon$-greedy as a computationally efficient but more myopic exploration strategy – this has been successful in practice [48, 9] and also enjoys theoretical guarantees under assumptions about the easiness of exploration [18, 72]. Another approach is to assume access to an offline dataset that already has good coverage and so strategic exploration is no longer necessary. This setting is called hybrid RL [55] and we revisit this in Sec. 4.7.
@@
and also set $\mathcal{E}^{\operatorname{RL}}_{\operatorname{sq}}=\sum_{h=1}^{H}\mathcal{E}^{\operatorname{sq}}_{h}$. We first establish an optimism lemma for $f^{k}$.
Lemma 7.
Let $\ell=\ell_{\operatorname{sq}}$ and $\mathcal{D}_{h}$ be a dataset where the $i$-th datapoint is collected from $\pi^{i}$, and denote $\mathcal{D}=(\mathcal{D}_{1},\dots,\mathcal{D}_{H})$. Then under BC (Assump. 3), for any $\delta\in(0,1)$, let $\beta=2\ln(H|\mathcal{F}|/\delta)$ and define
@@
W.p.a.l. $1-\delta$, we have (a) $\sum_{i=1}^{n}\mathcal{E}^{\operatorname{RL}}_{\operatorname{sq}}(\pi^{i};\hat{f}^{\operatorname{op}})\leq 2H\beta$, and (b) $\min_{a}\hat{f}^{\operatorname{op}}_{1}(x_{1},a)\leq V^{\star}$.
Proof of Lem. 7.
By standard martingale concentration via Freedman’s inequality, w.p.a.l. $1-\delta$, for all $f,h$, we have
$\textstyle\sum_{i=1}^{n}\mathcal{E}^{\operatorname{sq}}_{h}(f,\pi^{i})$ $\textstyle\leq\ln(H|\mathcal{F}|/\delta)+L_{h}^{\operatorname{sq}}(f_{h},f_{h+1},\mathcal{D}_{h})$
$\textstyle-L_{h}^{\operatorname{sq}}(\mathcal{T}_{h}f_{h+1},f_{h+1},\mathcal{D}_{h}).$ (15)
Let $g^{f}_{h}\in\mathop{\rm arg\,min}_{g_{h}\in\mathcal{G}_{h}}L_{h}^{\star}(g_{h},f_{h+1},\mathcal{D}_{h})$ denote the empirical risk minimizer, as used in the definition of $\mathcal{C}_{\beta}^{\star}(\mathcal{D})$ (Eq. 13). Under the BC premise,
$\textstyle\sum_{i=1}^{n}\mathcal{E}^{\operatorname{sq}}_{h}(f,\pi^{i})$ $\textstyle\leq\ln(H|\mathcal{F}|/\delta)+L_{h}^{\operatorname{sq}}(f_{h},f_{h+1},\mathcal{D}_{h})$
$\textstyle-L_{h}^{\operatorname{sq}}(g^{f}_{h},f_{h+1},\mathcal{D}_{h}).$
Thus, any $f\in\mathcal{C}_{\beta}^{\operatorname{sq}}(\mathcal{D})$ satisfies $\sum_{i=1}^{n}\mathcal{E}^{\operatorname{sq}}_{h}(f,\pi^{i})\leq 2\beta$, which proves Claim (a). For Claim (b), we prove that $Q^{\star}\in\mathcal{C}_{\beta}^{\operatorname{sq}}(\mathcal{D})$: by Eq. 15 and non-negativity of $\mathcal{E}^{\operatorname{sq}}$, we have $L_{h}^{\operatorname{sq}}(\mathcal{T}f_{h+1},f_{h+1},\mathcal{D}_{h})-L_{h}^{\operatorname{sq}}(g^{f}_{h},f_{h+1},\mathcal{D}_{h})\leq\ln(H|\mathcal{F}|/\delta)=\beta$. Then, setting $f=Q^{\star}$ and applying $Q^{\star}_{h}=\mathcal{T}^{\star}_{h}Q^{\star}_{h+1}$ shows that $Q^{\star}$ satisfies the version space condition. Thus, $Q^{\star}\in\mathcal{C}_{\beta}^{\operatorname{sq}}(\mathcal{D})$ and Claim (b) follows by definition of $\hat{f}^{\operatorname{op}}$. ∎
By Lem. 7, we have $\sum_{k=1}^{K}V^{\pi^{k}}-V^{\star}\leq\sum_{k=1}^{K}V^{\pi^{k}}-\min_{a}f^{k}_{1}(x_{1},a)$, which can be further decomposed by the performance difference lemma (PDL) [2, 35].
@@
The final step is to bound $\sum_{k=1}^{K}\mathcal{E}_{\operatorname{sq}}^{\operatorname{RL}}(f^{k},\pi^{k})$. By Lem. 7, we have that $\sum_{i<k}\mathcal{E}_{\operatorname{sq}}^{\operatorname{RL}}(f^{k},\pi^{i})\lesssim H\beta$ for all $k$, which is very similar except that the expectations are taken under previous policies $\pi^{<k}$ instead of $\pi^{k}$. It turns out that the eluder dimension can establish a link between the two, by using the following ‘pigeonhole principle’ lemma:
@@
Fix a number $N\in\mathbb{N}$, a sequence of functions $\psi^{(1)},\dots,\psi^{(N)}\in\Psi$, and a sequence of distributions $p^{(1)},\dots,p^{(N)}\in\mathcal{P}$. If for all $j\in[N]$, $\sum_{i<j}|\mathbb{E}_{p^{(i)}}\psi^{(j)}|^{q}\leq\beta^{q}$, then we have $\sum_{j=1}^{N}\lvert\mathbb{E}_{p^{(j)}}\psi^{(j)}\rvert\leq 2\operatorname{EluDim}_{q}(\Psi,\mathcal{P},N^{-1})\cdot(E+\beta^{q}\ln(EN))$, where $E:=\sup_{p\in\mathcal{P},\psi\in\Psi}\left\lvert\mathbb{E}_{p}\psi\right\rvert$ is the envelope.
Interpreting $\psi^{(i)}$ as the regression error at round $i$, Lem. 9 essentially states that ratio of (online) out-of-distribution errors (i.e., $\psi^{(i)}$ measured under $p^{(i)}$) to the (offline) in-distribution errors (i.e., $\psi^{(i)}$ measured under $p^{(1)},\dots,p^{(i-1)}$) is bounded by the eluder dimension. This lemma generalizes [54, 33, 42, 61] and we provide its proof in the appendix as Lem. 19.
Going back to the regret decomposition of Eq. 16, the pigeonhole lemma implies that $\sum_{k=1}^{K}\mathcal{E}^{RL}_{\operatorname{sq}}(f^{k},\pi^{k})\leq\widetilde{\mathcal{O}}(d_{\operatorname{sq}}H\beta)$. Thus, we have shown the desired regret bound $\widetilde{\mathcal{O}}(H\sqrt{Kd_{\operatorname{sq}}\beta})$.
If UA is true, we perform a change of measure to the uniform action distribution: $\sum_{k=1}^{K}\mathcal{E}^{RL}_{\operatorname{sq}}(f^{k},\pi^{k})\leq A\sum_{k=1}^{K}\sum_{h=1}^{H}\mathbb{E}_{\pi^{k}\circ_{h}\operatorname{unif}(\mathcal{A})}((f^{h}-\mathcal{T}_{h}f^{k}_{h+1})(x_{h},a_{h}))^{2}\lesssim Ad_{\operatorname{sq}}^{\operatorname{V}}H\beta$. Plugging into Eq. 16 gives the desired PAC bound $\widetilde{\mathcal{O}}(H\sqrt{AKd_{\operatorname{sq}}^{\operatorname{V}}\beta})$. This finishes the proof of Thm. 7. ∎
@@ 4.3 Verifying Assumptions for Low-Rank MDPs
We now show that $\mathcal{F}^{\operatorname{lin}}$ satisfies BC (Assump. 3).
@@
In a low-rank MDP, under Assump. 4, $\mathcal{F}^{\operatorname{lin}}$ satisfies Bellman Completeness (Assumps. 3 and 6).
Proof of Lem. 10.
@@
Moreover, we can also show that the V-type eluder dimension is bounded by the rank $d$ of the low-rank MDP, as defined in Def. 2.
@@
In a low-rank MDP with rank $d$, we have $\operatorname{EluDim}_{1}(\Psi^{\operatorname{V}}_{h},\mathcal{D}^{\operatorname{V}}_{h},\varepsilon)\leq\mathcal{O}(d\ln(d/\varepsilon))$ for all steps $h\in[H]$ and function classes $\Psi^{\operatorname{V}}_{h}\subset\mathcal{X}\to\mathbb{R}$.
@@
This can be proved by applying an elliptical potential argument to the decomposition in Eq. 17; for example, see Theorem G.4 of [61]. ∎
@@
4.4 First-Order Bounds for Online RL with Optimistic BCE Regression
As we learned from the CSC warmup, algorithms with the squared loss can be sub-optimal in small-cost problems. We also learned that simply swapping the loss function for the bce loss can yield first-order bounds that are more adaptive and sample efficient. We now show that this observation smoothly extends to RL as well.
In this subsection, we analyze Alg. 2 with the bce loss $\ell_{\operatorname{bce}}$ and derive improved first-order bounds. The intuition is that the Cauchy-Schwarz step in the proof of Thm. 7, while tight in the worst-case, is rather loose in many benign problems. We improve that step by leveraging Eq. 2 from the CSC warmup.
Before stating guarantees with the bce loss, we first define the eluder dimension which measures discrepancy with the Bernoulli squared hellinger distance. Let $\Psi_{h}^{\operatorname{bce}}=\{\delta^{\operatorname{Ber}}_{h}(\cdot;f):f\in\mathcal{F}\}$ where

### v3

@@ 4 Setup for Reinforcement Learning
In the rest of this paper, we will see that these observations and insights generally transfer to more complex decision making setups, in particular reinforcement learning (RL). Compared to the CSC setting, two new challenges of RL are that (1) the learner receives feedback only for the chosen action (a.k.a., partial or bandit feedback) and (2) the learner sequentially interacts with the environment over multiple time steps. As before, we focus on value-based algorithms with function approximation and prove bounds for problems with high-dimensional observations, i.e., beyond the finite tabular setting.
@@ 4.1 Problem Setup
5 Online RL with Squared-Loss Regression
Algorithm 1 Policy Data Collection
@@ 5 Online RL with Squared-Loss Regression
BC ensures that the TD-style regression which bootstraps on the next prediction is realizable, playing the same role as realizability (Asm. 1) in the CSC setting. In fact, BC implies realizability in $Q^{\star}$: $Q_{h}^{\star}\in\mathcal{F}_{h}$ for all $h$, which can be verified by using the Bellman optimality equations and induction from $h=H\to 1$. While appealing, $Q^{\star}$-realizability is not sufficient for sample efficient RL [62, 28, 71] and TD learning can diverge or converge to bad points with realizability alone [59, 53, 39]. We note that $Q^{\star}$-realizability becomes sufficient when combined with low coverability and generative access to the MDP [48], where the learner can reset to any previously observed states. We believe that the techniques in this paper can lead to first- and second-order bounds with realizability plus generative access for example. However, we do not pursue this direction here since exchanging BC for other conditions is orthogonal to our study of loss functions.
@@
Fix any set $\mathcal{S}$, function class $\Psi=\{\psi:\mathcal{S}\to{\mathbb{R}}\}$, distribution class =$\mathcal{M}=\{\mu:\Delta(\mathcal{S})\}$, threshold $\varepsilon_{0}$, and number $q\in\mathbb{N}$. The $\ell_{q}$-eluder dimension $\operatorname{EluDim}_{q}(\Psi,{\mathcal{M}},\varepsilon_{0})$ is defined as the length of the longest sequence $\mu^{(1)},\dots,\mu^{(L)}\subset\mathcal{M}$ s.t. $\exists\varepsilon\geq\varepsilon_{0}$, $\forall t\in[L]$, $\exists\psi\in\Psi$ s.t. $\left\lvert\mathbb{E}_{\mu^{(t)}}[\psi]\right\rvert>\varepsilon$ but $\sum_{i<t}\left\lvert\mathbb{E}_{\mu^{(i)}}[\psi]\right\rvert^{q}\leq\varepsilon^{q}$.
@@
The distribution class $\mathcal{M}$ will be the set of all visitation distributions by any policy, i.e., $\mathcal{M}_{h}=\{x,a\mapsto d^{\pi}_{h}(x,a):\pi\in\Pi\}$ where $d^{\pi}_{h}(x,a)$ is the state-action visitation distribution of $\pi$ at time step $h$. If UA is enabled, then we will have $\mathcal{S}=\mathcal{X}$ and the “V-type” distribution class is $\mathcal{M}^{\operatorname{V}}_{h}=\{x\mapsto d^{\pi}_{h}(x):\pi\in\Pi\}$ where $d^{\pi}_{h}(x)$ is the state visitation distribution of $\pi$ at time step $h$. Moreover we let $\Psi_{h}^{\operatorname{sq},\operatorname{V}}=\{\mathbb{E}_{a\sim\operatorname{unif}(\mathcal{A})}[\psi(x,a)]:\psi\in\Psi^{\operatorname{sq}}_{h}\}$ denote the “V-type” function class, where “V-type” refers to the fact that the functions only depend on state and not action. Thus, define the eluder dimension for squared loss:
@@
We now state the guarantees for Alg. 2 with the squared loss $\ell_{\operatorname{sq}}$, which recovers the main results of Jin, Liu and Miryoosefi [35].
@@
Under Asm. 3, for any $\delta\in(0,1)$, w.p.a.l. $1-\delta$, Alg. 2 with the squared loss $\ell_{\operatorname{sq}}$ and $\beta=2\ln(H|\mathcal{F}|/\delta)$ enjoys the following:
@@
Computational complexity of version space algorithms. The algorithms we present for online and offline RL optimize over version spaces to establish optimism and pessimism, and this optimization over version space is computationally inefficient in general [18]. The computational hardness comes from the non-convex optimization, not necessarily from any loss function itself. However, we note that the version space optimization is oracle-efficient in the one-step $H=1$ setting, a.k.a. contextual bandits [25, 22, 65]. In the RL setting, there are also approaches to mitigate the computational hardness of optimism. One common approach in practice is to use myopic exploration strategies such as epsilon-greedy [51, 10], which also has bounded regret when exploration is easy [19, 76]. Another approach which we later present is to work in the hybrid RL setting, where the learner can both interact with the MDP in an online manner and has access to offline data with good coverage [58]. In Sec. 8, we will see that our results for all loss functions naturally carry over to the hybrid setting, giving us both computational and statistical efficiency.
We now prove the main theorem for squared loss Thm. 7. We prove the nested lemmas in the Appendix.
@@
and also set $\mathcal{E}^{\operatorname{RL}}_{\operatorname{sq}}=\sum_{h=1}^{H}\mathcal{E}^{\operatorname{sq}}_{h}$. We first establish an optimism lemma for $f^{k}$ and prove it in the appendix. {restatable}lemmaSqRLOptimismLemma Let $\ell=\ell_{\operatorname{sq}}$ and $\mathcal{D}_{h}$ be a dataset where the $i$-th datapoint is collected from $\pi^{i}$, and denote $\mathcal{D}=(\mathcal{D}_{1},\dots,\mathcal{D}_{H})$. Then under BC (Asm. 3), for any $\delta\in(0,1)$, let $\beta=2\ln(H|\mathcal{F}|/\delta)$ and define
@@
W.p.a.l. $1-\delta$, we have (a) $\sum_{i=1}^{n}\mathcal{E}^{\operatorname{RL}}_{\operatorname{sq}}(\pi^{i};\hat{f}^{\operatorname{op}})\leq 2H\beta$, and (b) $\min_{a}\hat{f}^{\operatorname{op}}_{1}(x_{1},a)\leq V^{\star}$. By Sec. 5, we have $\sum_{k=1}^{K}V^{\pi^{k}}-V^{\star}\leq\sum_{k=1}^{K}V^{\pi^{k}}-\min_{a}f^{k}_{1}(x_{1},a)$, which can be further decomposed by the performance difference lemma (PDL) [2, 37].
@@
The final step is to bound $\sum_{k=1}^{K}\mathcal{E}_{\operatorname{sq}}^{\operatorname{RL}}(f^{k},\pi^{k})$. By Sec. 5, we have that $\sum_{i<k}\mathcal{E}_{\operatorname{sq}}^{\operatorname{RL}}(f^{k},\pi^{i})\lesssim H\beta$ for all $k$, which is very similar except that the expectations are taken under previous policies $\pi^{<k}$ instead of $\pi^{k}$. It turns out that the eluder dimension can establish a link between the two, via the following “pigeonhole principle” lemma:
@@
Fix a number $N\in\mathbb{N}$, a sequence of functions $\psi^{(1)},\dots,\psi^{(N)}\in\Psi$, and distributions $\mu^{(1)},\dots,\mu^{(N)}\in\mathcal{M}$. Suppose $\sum_{i<j}|\mathbb{E}_{\mu^{(i)}}[\psi^{(j)}]|^{q}\leq\beta^{q}$ for all $j\in[N]$. Then we have $\sum_{j=1}^{N}\lvert\mathbb{E}_{\mu^{(j)}}[\psi^{(j)}]\rvert\leq 2\operatorname{EluDim}_{q}(\Psi,\mathcal{M},N^{-1})\cdot(E+\beta^{q}\ln(EN))$, where $E:=\sup_{\mu\in\mathcal{M},\psi\in\Psi}\left\lvert\mathbb{E}_{\mu}[\psi]\right\rvert$ is the envelope.
If we interpret $\psi^{(i)}$ as the regression error at round $i$, then Lem. 8 essentially states that ratio of (online) out-of-distribution errors (i.e., $\psi^{(i)}$ measured under $p^{(i)}$) to the (offline) in-distribution errors (i.e., $\psi^{(i)}$ measured under $p^{(1)},\dots,p^{(i-1)}$) is bounded by the eluder dimension. This generalizes similar results from [57, 35, 44, 64, 74].
Finally, going back to the regret decomposition in Eq. 15, applying the pigeonhole lemma implies that $\sum_{k=1}^{K}\mathcal{E}^{RL}_{\operatorname{sq}}(f^{k},\pi^{k})\leq\widetilde{\mathcal{O}}(d_{\operatorname{sq}}H\beta)$. Thus, we have shown that $\sum_{k=1}^{K}V^{\pi^{k}}-f^{k}_{1}(x_{1},\pi^{k}(x_{1}))\leq\widetilde{\mathcal{O}}(H\sqrt{Kd_{\operatorname{sq}}\beta})$, which proves the desired regret bound.
Moreover, if the uniform action (UA) flag is set, we also perform a change of measure so that: $\sum_{k=1}^{K}\mathcal{E}^{RL}_{\operatorname{sq}}(f^{k},\pi^{k})\leq A\sum_{k=1}^{K}\sum_{h=1}^{H}\mathbb{E}_{\pi^{k}\circ_{h}\operatorname{unif}(\mathcal{A})}((f^{h}-\mathcal{T}_{h}f^{k}_{h+1})(x_{h},a_{h}))^{2}\lesssim Ad_{\operatorname{sq}}^{\operatorname{V}}H\beta$. Plugging into Eq. 15 proves the desired PAC bound of $\widetilde{\mathcal{O}}(H\sqrt{AKd_{\operatorname{sq}}^{\operatorname{V}}\beta})$. This finishes the proof. ∎
@@ 5.1 Verifying Assumptions for Low-Rank MDPs
We now show that $\mathcal{F}^{\operatorname{lin}}$ satisfies BC (Asm. 3).
@@
In a low-rank MDP, under Asm. 4, $\mathcal{F}^{\operatorname{lin}}$ satisfies Bellman Completeness (Asms. 3 and 6).
Proof of Lem. 9.
@@
Moreover, we can also show that the V-type eluder dimension is bounded by the rank $d$ of the low-rank MDP, as defined in Def. 2. We note this applies to both $\ell_{1}$ and $\ell_{2}$ eluder dimensions.
@@
In a low-rank MDP with rank $d$, we have $\operatorname{EluDim}_{1}(\Psi^{\operatorname{V}}_{h},\mathcal{M}^{\operatorname{V}}_{h},\varepsilon)\leq\operatorname{EluDim}_{2}(\Psi^{\operatorname{V}}_{h},\mathcal{M}^{\operatorname{V}}_{h},\varepsilon)\leq\mathcal{O}(d\ln(d/\varepsilon))$ for all steps $h\in[H]$ and function classes $\Psi^{\operatorname{V}}_{h}\subset\mathcal{X}\to\mathbb{R}$.
@@
$\operatorname{EluDim}_{2}(\Psi^{\operatorname{V}}_{h},\mathcal{M}^{\operatorname{V}}_{h},\varepsilon)\leq\mathcal{O}(d\ln(d/\varepsilon))$ can be proved by applying a standard elliptical potential argument to the decomposition in Eq. 16; for example, see [35] or Theorem G.4 of [64]. Then, $\operatorname{EluDim}_{1}(\Psi^{\operatorname{V}}_{h},\mathcal{M}^{\operatorname{V}}_{h},\varepsilon)\leq\operatorname{EluDim}_{2}(\Psi^{\operatorname{V}}_{h},\mathcal{M}^{\operatorname{V}}_{h},\varepsilon)$ is a simple consequence of the fact that $\sqrt{\sum_{i}x_{i}^{2}}\leq\sum_{i}|x_{i}|$ [64, Lemma 5.4]. ∎
@@
6 First/Second-Order Bounds for Online RL
As we learned from the CSC warmup, algorithms with the squared loss can be sub-optimal in small-cost or small-variance problems. The intuition is that squared loss regression bounds do not capture the underlying heteroskedastic variance; indeed, minimizing squared loss implicitly assumes that the underlying distribution is a homoskedastic Gaussian. We also learned that simply swapping the loss function for the bce loss or mle loss can yield first-order or second-order bounds that are more sample efficient in small-loss or small-variance settings. We now show that this observation smoothly extends to RL as well. In particular, we will sharpen the Cauchy-Schwarz step in the proof of Thm. 7 by leveraging Eq. 2 from the CSC warmup.
6.1 First-Order Bounds with BCE Regression
In this subsection, we analyze Alg. 2 with the bce loss $\ell_{\operatorname{bce}}$ and derive improved first-order bounds. Before stating guarantees with the bce loss, we first define the eluder dimension which measures discrepancy with the Bernoulli squared hellinger distance. Let $\Psi_{h}^{\operatorname{bce}}=\{\delta^{\operatorname{Ber}}_{h}(\cdot;f):f\in\mathcal{F}\}$ where